OpenStudy (anonymous):

WILL GIVE MEDAL + FAN Which of the following is the root of the equation 4x^2+2x^3+kx+7=0, where k is an integer. A.2 B.4 C.7/2 D.2/7

OpenStudy (anonymous):

2x^3+4x^2+kx+7=0 let roots are a,b,c sum of roots =a+b+c=-4/2=-2 product of roots=-7/4

OpenStudy (anonymous):

abd=-7/4 ab+bc+ca=k/4

OpenStudy (anonymous):

sorry abc=-7/4

OpenStudy (anonymous):

Ok the answers is simply to use the Rational Zero Theorem (: